Find the ratio at which 50% alcohol be mixed with 75% alcohol to form a
solution of 60% alcohol

Answer by ewatrrr(24785)


Hi,
using the example of forming a 100L solution
Let x and (100-x)represent the amount of the 75% and 50% solutions respectively
Mixing to the form 100L of the 60% solution
.75x + .50(100L-x) = .60*100L
.25x = .10*100L
x = 10L/.25 = 40L, the 75%, 60L, the 50%
Ratio of the 50% alcohol be mixed with 75% alcohol
60/40 = 6/4 = 3/2 OR 3:2